# On the existential theory of the completions of a global field

### v2

@@ 1. Introduction
The (common) $\mathfrak{L}_{\rm val}$-theory of all $\mathbb{Q}_{p}$ and the $\mathfrak{L}_{\rm val}$-theory of almost all $\mathbb{Q}_{p}$ is decidable.
@@
Let $K$ be a number field. The $\mathfrak{L}_{\rm ring}$-theory of all completions of $K$ and the $\mathfrak{L}_{\rm ring}$-theory of almost all completions of $K$ is decidable. In particular, it is decidable whether a given polynomial over $K$ has a zero in all respectively almost all completions of $K$ [^1] For a sketch see Remark 6.8.
@@
For completions of global function fields, there is at this point little hope to decide the common theory of all completions, since even the theory of a single completion $\mathbb{F}_{q}(\!(t)\!)$ seems out of reach, in spite of progress such as [Kuh01]. For the existential theory (that is, essentially, deciding which polynomials have zeros), the following result of Anscombe and the second author is somewhat orthogonal:
Theorem 1.3 ([AF23]).
The existential $\mathfrak{L}_{\rm val}$-theory of all respectively almost all $\mathbb{F}_{p}(\!(t)\!)$, where $p$ runs over the prime numbers, is decidable.
@@ 2. $(f)$
In fact, we will not only prove decidability but also axiomatise these theories, where possible for general function fields of one variable. Moreover we will prove the results not just for existential sentences but for universal/existential sentences, i.e. finite boolean combinations of existential sentences (so that in particular we can decide not only whether a given polynomial has a zero in all respectively all but finitely many completions, but also whether it has a zero in at most finitely many respectively no completions).
@@ 2. Notation and preliminaries
We fix a field $C$ and denote by $T_{0}$ the $\mathfrak{L}_{\rm val}(C)$-theory of valued fields extending the trivially valued field $C$. For an $\mathfrak{L}_{\rm val}(C)$-theory $T\supseteq T_{0}$ and an $\mathfrak{L}_{\rm ring}(C)$-theory $R$ we denote by $T(R)$ the theory of $(F,v)\models T$ with $Fv\models R$. Here we view $C$ as embedded into $Fv$ via the residue map.
@@
($\ast$) ${\rm Th}_{\forall/\exists}(F,v)={\rm Th}_{\forall/\exists}(F^{\prime},v^{\prime})\mbox{ whenever }(F,v),(F^{\prime},v^{\prime})\models T\mbox{ with }{\rm Th}_{\forall/\exists}(Fv)={\rm Th}_{\forall/\exists}(F^{\prime}v^{\prime}).$
@@ 3. $(c)$
Moreover, if we weaken the conclusion of ()(*)( ∗ ) to ${\rm Th}_{\forall/\exists}(F)={\rm Th}_{\forall/\exists}(F^{\prime})$, then (a), (b) and (c) hold if we interpret each “$\forall$/$\exists$ / ∃” to refer only to universal/existential $\mathfrak{L}_{\rm ring}(C)$-sentences.
@@
This can be seen directly, or deduced from Corollary 2.23 of [AF23]: The universal/existential sentences of $\mathfrak{L}_{1}=\mathfrak{L}_{\rm ring}(C)$ and $\mathfrak{L}_{2}=\mathfrak{L}_{\rm val}(C)$ form fragments in the terminology of that paper, and together with the theories $R_{0}$ respectively $T$ and the map $\sigma\colon(F,v)\mapsto Fv$, they form a bridge $B$. With $\hat{B}$ the same bridge but now all $\mathfrak{L}_{\rm ring}(C)$- respectively $\mathfrak{L}_{\rm val}(C)$-sentences, and $\iota$ the map on sentences obtained from interpreting the residue field in the valued field, we obtain an arch $A=(B,\hat{B},\iota)$. Assumption (i) of that corollary is satisfied by our hypotheses on $C$ and $T$, assumption (ii) is satisfied by the definition of $R_{0}$, and assumption (iii) is satisfied by ($\ast$2.1). Part (a) of our lemma follows from (III) of the corollary, part (b) follows from (II), and part (c) follows from (I) together with [ADF23, Lemma 2.24]. The “moreover” part follows just the same, after changing $\mathfrak{L}_{2}$ to $\mathfrak{L}_{\rm ring}(C)$. ∎
@@ 3. Existential $\mathfrak{L}_{\rm val}$-theory without parameters
We start by studying the universal/existential $\mathfrak{L}_{\rm val}$-theory of all and almost all completions of a function field without any parameters. Here our background theory is the $\mathfrak{L}_{\rm val}$-theory $T_{1}$ of equicharacteristic henselian nontrivially valued fields of fixed characteristic $p\geq 0$, which for technical reasons we view as an $\mathfrak{L}_{\rm val}(C)$-theory for $C=\mathbb{F}_{p}$ respectively $C=\mathbb{Q}$.
@@
By [AF16, Thm. 1.1], $(F,v)$ and $(F^{\prime},v^{\prime})$ have the same existential $\mathfrak{L}_{\rm val}$-theory, and therefore also the same universal/existential $\mathfrak{L}_{\rm val}$-theory. ∎
If $K/k$ is a function field and $v\in\mathbb{P}_{K}$, we denote by $(\hat{K}_{v},\hat{v})$ the completion of $(K,v)$, which is an equicharacteristic henselian nontrivially valued field with residue field $\hat{K}_{v}\hat{v}=Kv$.
@@
As ($\ast$2.1) holds for $T=T_{1}$ by Proposition 3.1, and $(\hat{K}_{v},\hat{v})$ is a model of $T_{1}$, Lemma 2.1(a) gives that ${\rm Th}_{\forall/\exists}(\hat{K}_{v},\hat{v})=T_{1}({\rm Th}(Kv))_{\forall/\exists}$ for every $v\in\mathbb{P}_{K}$. Lemma 2.1(c) shows that for each $P\subseteq\mathbb{P}_{K}$,
@@
We now turn to the case where $k=\mathbb{F}_{q}$ is finite. Then $Kv$ is a finite field for each $v\in\mathbb{P}_{K}$, and the completion $(\hat{K}_{v},\hat{v})$ of $(K,v)$ is isomorphic to the Laurent series field $(Kv(\!(t)\!),v_{t})$ with the $t$-adic valuation $v_{t}$ [Ser79, Thm. II.2]. For $f\in k[t,x]$ irreducible we denote by $K_{f}$ the global function field ${\rm Quot}(k[t,x]/(f))$.
@@
Every $Kv$ is a finite extension of the field of constants $k$, and so $|Kv|=q^{[Kv:k]}$. Without loss of generality, $d:={\rm deg}(f)>1$. The genus $g$ of $K$ is at most $(d-1)^{2}/2$ [Sti09, Cor. 3.11.4], and the Hasse–Weil bound implies that $K$ has a place of degree $n$ whenever $n\geq 4g+3$ [Sti09, Cor. 5.2.10(c)]. This holds in case $n>2d^{2}$. ∎
@@
[…]
@@ 4. Existential $\mathfrak{L}_{\rm val}$-theory with constant parameters, and the perfect hull
Next we consider two cases where adding constants from a subfield $C$ of $K$ does not do much harm: This essentially happens when the fields involved are perfect. Our background theory now is the $\mathfrak{L}_{\rm val}(C)$-theory $T_{2}\supseteq T_{1}$ of henselian nontrivially valued fields $(F,v)$ extending the trivially valued field $C$ with $Fv$ perfect and $Fv/C$ separable (which implies $C$ perfect).
@@
Corollary 5.7 of [AF16] implies that ${\rm Th}_{\forall/\exists}(L,w)=T_{2}({\rm Th}(Lw))_{\forall/\exists}$ for every $(L,w)\models T_{2}$. We assume ${\rm Th}_{\forall/\exists}(Fv)={\rm Th}_{\forall/\exists}(F^{\prime}v^{\prime})$ and have to show that ${\rm Th}_{\forall/\exists}(F,v)={\rm Th}_{\forall/\exists}(F^{\prime},v^{\prime})$, and so we can replace $(F^{\prime},v^{\prime})$ by an elementary extension to assume without loss of generality that $F^{\prime}v^{\prime}$ is $|Fv|^{+}$-saturated. As in particular ${\rm Th}_{\exists}(Fv)\subseteq{\rm Th}_{\exists}(F^{\prime}v^{\prime})$ we obtain an embedding $Fv\rightarrow F^{\prime}v^{\prime}$ over $C$ [CK90, Lemma 5.2.1], which extends to an embedding $(Fv(\!(t)\!),v_{t})\rightarrow(F^{\prime}v^{\prime}(\!(t)\!),v_{t})$, proving
@@
Let $C=K^{\rm perf}$. By Łoś’s theorem [CK90, Thm. 4.1.9], the $\mathfrak{L}_{\rm val}(C)$-theory of almost all $(\hat{K}_{v})^{\rm perf}$ is the $\mathfrak{L}_{\rm val}(C)$-theory of the ultraproducts $(K_{\mathcal{U}},v_{\mathcal{U}})=\prod_{v\in\mathbb{P}_{K}}((\hat{K}_{v})^{\rm perf},\hat{v})/\mathcal{U}$, where $\mathcal{U}$ runs over the nonprincipal ultrafilters on $\mathbb{P}_{K}$. As every $((\hat{K}_{v})^{\rm perf},\hat{v})$ is a perfect henselian nontrivially valued field, so is $(K_{\mathcal{U}},v_{\mathcal{U}})$. In particular, $K_{\mathcal{U}}v_{\mathcal{U}}$ is perfect. Since every $c\in C^{\times}$ has nonzero value for at most finitely many $v\in\mathbb{P}_{K}$, $v_{\mathcal{U}}$ is trivial on $C$. Thus $(K_{\mathcal{U}},v_{\mathcal{U}})\models T_{2}$. Similarly, the $\mathfrak{L}_{\rm val}(C)$-theory of almost all $\hat{L}_{v}$ is the $\mathfrak{L}_{\rm val}(C)$-theory of the ultraproducts $(L_{\mathcal{U}},w_{\mathcal{U}})=\prod_{v\in\mathbb{P}_{K}}(\hat{L}_{v},\hat{v})/\mathcal{U}$, and also $(L_{\mathcal{U}},w_{\mathcal{U}})\models T_{2}$. Moreover, as $(\hat{K}_{v})^{\rm perf}\hat{v}=(Kv)^{\rm perf}=Lv$, we have that $K_{\mathcal{U}}v_{\mathcal{U}}=\prod_{v\in\mathbb{P}_{K}}(Kv)^{\rm perf}/\mathcal{U}=L_{\mathcal{U}}w_{\mathcal{U}}$. Proposition 4.1 and Lemma 2.1(a,c) therefore give that
@@
As noted in the introduction, we could always interpret “almost all” in the sense of Dirichlet density $1$ (as defined for example in [FJ08, Chapter 6.3] or [Ros02, p. 123]). Indeed, for Lemma 4.4, this is explicit in its proof. This then carries over to Theorem 3.4 if in the proof of Proposition 3.2 one lets $P_{0}$ run over sets of Dirichlet density 00 rather than only over finite sets. It also carries over to Theorem 4.5 (and similarly the theorems in the following two sections) if in the proof of Proposition 4.3 one lets $\mathcal{U}$ run only over those ultrafilters on $\mathbb{P}_{K}$ that contain the filter of subsets of Dirichlet density $1$ rather than over all nonprincipal ultrafilters.
@@ 5. Existential $\mathfrak{L}_{\rm ring}$-theory with parameters
Let $C$ be a field of imperfect exponent at most $1$. If $(F,v)$ and $(F^{\prime},v^{\prime})$ are henselian nontrivially valued separable extensions of the trivially valued field $C$ such that $Fv$ and $F^{\prime}v^{\prime}$ have the same universal/existential $\mathfrak{L}_{\rm ring}(C)$-theory and every geometrically integral smooth affine $C$-variety has a rational point over both $Fv$ and $F^{\prime}v^{\prime}$, then $F$ and $F^{\prime}$ have the same universal/existential $\mathfrak{L}_{\rm ring}(C)$-theory, i.e. ($\ast$2.1) holds for $T=T_{3}$ with the weaker conclusion for $\mathfrak{L}_{\rm ring}(C)$ instead of $\mathfrak{L}_{\rm val}(C)$.
@@
We claim that by enlarging $C$ we can assume without loss of generality that $C$ is separably closed in $Fv$ and $F^{\prime}v^{\prime}$. Indeed, if $D$ is the separable algebraic closure of $C$ in $Fv$, then $C$ and $D$ have the same imperfect exponent [FJ08, Lemma 2.7.3]. Moreover, since $Fv$ and $F^{\prime}v^{\prime}$ have the same universal/existential $\mathfrak{L}_{\rm ring}(K)$-theory, the relative algebraic closure of $C$ in $Fv$ is isomorphic to the one in $F^{\prime}v^{\prime}$ [FJ08, Lemma 20.6.3(b)], and so we can choose an isomorphism of $D$ with the separable algebraic closure of $C$ in $F^{\prime}v^{\prime}$. As $v$ is trivial on $C$, we can identify the separable algebraic closure of $C$ in $F$ with a subfield of $D$ via the residue map, and since $v$ is henselian, this subfield is actually all of $D$ (cf. [AF16, Lemma 2.3]). Similarly we identify $D$ with the separable algebraic closure of $C$ in $F^{\prime}$. Note that $F$ and $F^{\prime}$ remain separable over $D$. Finally, every geometrically integral smooth affine $D$-variety has a rational point over $Fv$ and $F^{\prime}v^{\prime}$: Such a $D$-variety is the base change of a geometrically integral smooth affine $C^{\prime}$-variety $V$ for a finite subextension $C^{\prime}/C$ of $D/C$. The Weil restriction $W={\rm res}_{C^{\prime}/C}(V)$ is a geometrically integral smooth affine $C$-variety (see [Sch94, Prop. 4.9] and [Die01, Lemma 1.10]), hence $W(Fv)\neq\emptyset$ by assumption. From the morphism $W_{C^{\prime}}\rightarrow V$ (see [Sch94, 4.2.5]) we conclude that also $V(Fv)\neq\emptyset$, and similarly $V(F^{\prime}v^{\prime})\neq\emptyset$.
So from now on assume that $C$ is separably closed in $Fv$ and $F^{\prime}v^{\prime}$. We already argued that then it is also separably closed in $F$. Since $F/C$ is separable, $C$ is therefore relatively algebraically closed in $F$. Together with the assumption of imperfect exponent at most $1$ this implies that $F/C$ is regular [FJ08, Lemma 2.7.5]. Similarly, $F^{\prime}/C$ is regular. Therefore, by [Dit23, Lemma 4.2], $F$ and $F^{\prime}$ have the same existential $\mathfrak{L}_{\rm ring}(C)$-theory, and therefore the same universal/existential $\mathfrak{L}_{\rm ring}(C)$-theory, as soon as every geometrically integral smooth affine $C$-variety $V$ has an $F$-rational point if and only if it has an $F^{\prime}$-rational point. But every such $V$ has an $Fv$-rational point by assumption, and this $Fv$-rational point lifts to an $F$-rational point of $V$ since $v$ is henselian (see for instance [BLR90, §2.3 Prop. 5]), and similarly to an $F^{\prime}$-rational point. This concludes the proof. ∎
@@
Each $(\hat{K}_{v},\hat{v})$ is henselian nontrivially valued. Furthermore, each $\hat{K}_{v}$ is separable over $K$, since a uniformiser $\pi\in K$ for $v$ is a $p$-basis for both $K$ and $\hat{K}_{v}$. For each $x\in K^{\times}$, $v(x)\neq 0$ for only finitely many $v\in\mathbb{P}_{K}$, so models of the $\mathfrak{L}_{\rm val}(K)$-theory of almost all $\hat{K}_{v}$ have valuation trivial on $K$. Finally, such a model has pseudofinite residue field by Lemma 4.4. ∎
@@
where now $\forall$/$\exists$ / ∃ denotes universal/existential $\mathfrak{L}_{\rm ring}(K)$-consequences. As $\bigcap_{\mathcal{U}}{\rm Th}(K_{\mathcal{U}}v_{\mathcal{U}})=\bigcap_{\mathcal{U}}{\rm Th}(\prod_{v\in\mathbb{P}_{K}}Kv/\mathcal{U})$ is the $\mathfrak{L}_{\rm ring}(K)$-theory of almost all $Kv$, which by Lemma 4.4 is the $\mathfrak{L}_{\rm ring}(K)$-theory of pseudofinite fields containing $K$, the first claim follows. The decidability follows from Lemma 2.1(b) and Lemma 4.4. ∎

### v3

@@ 1. Introduction
The (common) $\mathfrak{L}_{\rm val}$-theory of all $\mathbb{Q}_{p}$ and the $\mathfrak{L}_{\rm val}$-theory of almost all $\mathbb{Q}_{p}$ are both decidable.
@@
Let $K$ be a number field. The $\mathfrak{L}_{\rm ring}$-theory of all completions of $K$ and the $\mathfrak{L}_{\rm ring}$-theory of almost all completions of $K$ are both decidable. In particular, it is decidable whether a given polynomial over $K$ has a zero in all respectively almost all completions of KK. [^1] For a sketch see Remark 6.9.
@@
For completions of global function fields, there is at this point little hope to decide the common theory of all completions, since even the theory of a single completion $\mathbb{F}_{q}(\!(t)\!)$ seems out of reach, in spite of progress such as [Kuh01]. For the existential theory (that is, essentially, deciding which polynomials have zeros), the following result of Anscombe and the second author (which can be obtained by combining Corollary 5.11(1c,d) and Proposition 5.9(1a) of [AF26]) is related but does not consider the family of completions of one fixed global function field like $\mathbb{F}_{p}(t)$, but instead varies $p$:
Theorem 1.3 ([AF26]).
The existential $\mathfrak{L}_{\rm val}$-theory of all $\mathbb{F}_{p}(\!(t)\!)$ and the existential $\mathfrak{L}_{\rm val}$-theory of almost all $\mathbb{F}_{p}(\!(t)\!)$, where $p$ runs over the prime numbers, are both decidable.
@@ 2. $(f)$
In fact, we will not only prove decidability but also axiomatise these theories, where possible for general function fields of one variable. Moreover we will prove the results not just for existential sentences but for universal/existential sentences, i.e. finite boolean combinations of existential sentences (so that in particular we can decide not only whether a given polynomial has a zero in all completions and whether it has a zero in almost all completions, but also whether it has a zero in at most finitely many completions, and whether it has a zero in no completion.
@@ 2. Notation and preliminaries
We fix a field $C$ and denote by $T_{0}$ the $\mathfrak{L}_{\rm val}(C)$-theory of valued fields extending the trivially valued field $C$. For an $\mathfrak{L}_{\rm val}(C)$-theory $T\supseteq T_{0}$ and an $\mathfrak{L}_{\rm ring}(C)$-theory $R$ we denote by $T(R)$ the theory of $(F,v)\models T$ with $Fv\models R$. Here we view $C$ as embedded into $Fv$ via the residue map. For example, in the following lemma, $T({\rm Th}_{\forall/\exists}(Fv))_{\forall/\exists}$ refers to the set of universal/existential consequences of the common $\mathfrak{L}_{\rm val}(C)$-theory of those $(L,w)\models T$ whose residue field $Lw$ satisfies the same universal/existential $\mathfrak{L}_{\rm ring}(C)$-sentences as $Fv$.
@@
[…]
@@ 3. $(c)$
Moreover, if we weaken the conclusion of ()(*) to ${\rm Th}_{\forall/\exists}(F)={\rm Th}_{\forall/\exists}(F^{\prime})$, then (a), (b) and (c) hold if we interpret each $\forall/\exists$” to refer only to universal/existential $\mathfrak{L}_{\rm ring}(C)$-sentences.
@@
This can be seen directly, or deduced from Corollary 2.23 of [AF26]: The universal/existential sentences of $\mathfrak{L}_{1}=\mathfrak{L}_{\rm ring}(C)$ and $\mathfrak{L}_{2}=\mathfrak{L}_{\rm val}(C)$ form fragments in the terminology of that paper, and together with the theories $R_{0}$ respectively $T$ and the map $\sigma\colon(F,v)\mapsto Fv$, they form a bridge $B$. With $\hat{B}$ the same bridge but now all $\mathfrak{L}_{\rm ring}(C)$- respectively $\mathfrak{L}_{\rm val}(C)$-sentences, and $\iota$ the map on sentences obtained from interpreting the residue field in the valued field, we obtain an arch $A=(B,\hat{B},\iota)$. Assumption (i) of that corollary is satisfied by our hypotheses on $C$ and $T$, assumption (ii) is satisfied by the definition of $R_{0}$, and assumption (iii) is satisfied by ($\ast$2.1). Part (a) of our lemma follows from (III) of the corollary, part (b) follows from (II), and part (c) follows from (I) together with [ADF23, Lemma 2.24]. The “moreover” part follows just the same, after changing $\mathfrak{L}_{2}$ to $\mathfrak{L}_{\rm ring}(C)$. ∎
For the convenience of the reader, we now list the theories $T\supseteq T_{0}$ which we will discuss in this work:
  1. $T_{1}$
    is the $\mathfrak{L}_{\rm val}$-theory of equicharacteristic henselian nontrivially valued fields of fixed characteristic $p\geq 0$, see Section 3.
  2. $T_{2}$
    is the $\mathfrak{L}_{\rm val}(C)$-theory of henselian nontrivially valued fields $(F,v)$ extending the trivially valued field $C$ with $Fv$ perfect and $Fv/C$ separable, see Section 4.
  3. $T_{3}$
    is the $\mathfrak{L}_{\rm val}(C)$-theory of henselian nontrivially valued fields $(F,v)$ extending the trivially valued field $C$ such that $F/C$ is separable and every geometrically integral smooth $C$-variety has $Fv$-rational points, see Section 5.
  4. $T_{4}$
    is the $\mathfrak{L}_{\rm val}(C)$-theory of henselian nontrivially valued fields $(F,v)$ extending the trivially valued field $C$ satisfying an additional axiom UR($t$), to be defined in Section 6.
@@ 3. Existential $\mathfrak{L}_{\rm val}$-theory without parameters
We start by studying the universal/existential $\mathfrak{L}_{\rm val}$-theory of all and almost all completions of a function field without any parameters. Here our background theory is the $\mathfrak{L}_{\rm val}$-theory $T_{1}$ of equicharacteristic henselian nontrivially valued fields of fixed characteristic $p\geq 0$, which for technical reasons we view as an $\mathfrak{L}_{\rm val}(C)$-theory for $C=\mathbb{F}_{p}$ respectively $C=\mathbb{Q}$, and the main result we will use about this theory is the following:
Theorem 3.1 ([AF16, Thm. 1.1]).
For every field $k$, the theory $T=T_{1}({\rm Th}_{\forall/\exists}(k))$ is $\exists$-complete, i.e. for every existential $\mathfrak{L}_{\rm val}$-sentence $\varphi$ we have either $T\models\varphi$ or $T\models\neg\varphi$.
@@
By Theorem 3.1, $(F,v)$ and $(F^{\prime},v^{\prime})$ have the same existential $\mathfrak{L}_{\rm val}$-theory, and therefore also the same universal/existential $\mathfrak{L}_{\rm val}$-theory. ∎
[…]
@@
[…]
@@
[…]
@@
Every KvKv is a finite extension of the field of constants $k$, and so $|Kv|=q^{[Kv:k]}$. Without loss of generality, $d:={\rm deg}(f)>1$. The genus $g$ of $K$ is at most $(d-1)^{2}/2$ [Sti09, Cor. 3.11.4], and the Hasse–Weil bound implies that $K$ has a place of degree $n$ whenever $n\geq 4g+3$ [Sti09, Cor. 5.2.10(c)]. This holds in case $n>2d^{2}$. ∎
@@
One could deduce Theorem 1.4(a,b) from the (a priori weaker) corresponding statement for $\mathfrak{L}_{\rm ring}$ instead of $\mathfrak{L}_{\rm val}$ if there were an existential $\mathfrak{L}_{\rm ring}$-formula $\varphi(x)$ and a universal $\mathfrak{L}_{\rm ring}$-formula $\psi(x)$ (all formulas here without parameters) such that both $\varphi$ and $\psi$ would define the valuation ring in all (respectively almost all) completions of $K$. However, since by Lemma 3.4 the smallest elementary class $\mathcal{F}\supseteq\{Kv:v\in\mathbb{P}_{K}\}$ contains pseudofinite fields extending $\overline{\mathbb{F}}_{q}$, [AF17, Thm. 5.1, Prop. 6.5(a), 6.20(a)] (with $C=\mathbb{Z}$, $\mathcal{K}=\{(\hat{K}_{v},\hat{v}):v\in\mathbb{P}_{K}\})$ proves that neither $\varphi$ nor $\psi$ can exist. We point out that although it is known that there exist $\exists\forall$-formulas and $\forall\exists$-formulas that define the valuation ring in all completions of $K$ (see [FP15, Thm. 1.4, 1.6], and [FJ17] for more on definable valuations), no universal/existential formula $\eta$ defines the valuation ring in almost all completions of $K$: Fix $v_{0}\in\mathbb{P}_{K}$ such that $q^{d}\in R(K)$ for every $q^{d}\geq|Kv_{0}|$ and such that $\eta$ defines the valuation ring in $\hat{K}_{v}$ for every $v$ with $|Kv|\geq|Kv_{0}|$. Then $\eta$ defines the valuation ring of the unique extension of $v_{0}$ in every finite extension of $\hat{K}_{v_{0}}$. Since both $\eta$ and $\neg\eta$ are logically equivalent to $\forall\exists$-formulas, and therefore go up chains [CK90, Exercise 3.1.8], this implies that also in the algebraic closure of $\hat{K}_{v_{0}}$, $\eta$ defines the valuation ring of the unique extension of $v_{0}$, contradicting quantifier elimination for algebraically closed fields.
@@ 4. Existential $\mathfrak{L}_{\rm val}$-theory with constant parameters, and the perfect hull
Next we consider two cases where adding constants from a subfield $C$ of $K$ does not do much harm: This essentially happens when the fields involved are perfect. Our background theory now is the $\mathfrak{L}_{\rm val}(C)$-theory $T_{2}\supseteq T_{1}$ of henselian nontrivially valued fields $(F,v)$ extending the trivially valued field $C$ with $Fv$ perfect and $Fv/C$ separable (which implies $C$ perfect). We will use the following result about this theory:
Theorem 4.1 ([AF16, Corollary 5.7]).
For every separable extension $k/C$ with $k$ perfect, the theory $T=T_{2}({\rm Th}(k))$ is $\exists$-$C$-complete, i.e. for any existential $\mathfrak{L}_{\rm val}(C)$-sentence $\varphi$ we have either $T\models\varphi$ or $T\models\neg\varphi$.
@@
Theorem 4.1 implies that ${\rm Th}_{\forall/\exists}(L,w)=T_{2}({\rm Th}(Lw))_{\forall/\exists}$ for every $(L,w)\models T_{2}$. We assume ${\rm Th}_{\forall/\exists}(Fv)={\rm Th}_{\forall/\exists}(F^{\prime}v^{\prime})$ and have to show that ${\rm Th}_{\forall/\exists}(F,v)={\rm Th}_{\forall/\exists}(F^{\prime},v^{\prime})$, and so we can replace $(F^{\prime},v^{\prime})$ by an elementary extension to assume without loss of generality that $F^{\prime}v^{\prime}$ is $|Fv|^{+}$-saturated. As in particular ${\rm Th}_{\exists}(Fv)\subseteq{\rm Th}_{\exists}(F^{\prime}v^{\prime})$ we obtain an embedding $Fv\rightarrow F^{\prime}v^{\prime}$ over $C$ [CK90, Lemma 5.2.1], which extends to an embedding $(Fv(\!(t)\!),v_{t})\rightarrow(F^{\prime}v^{\prime}(\!(t)\!),v_{t})$, proving
@@
[…]
@@
As noted in the introduction, we could always interpret “almost all” in the sense of Dirichlet density $1$ (as defined for example in [FJ08, Chapter 6.3] or [Ros02, p. 123]). Indeed, for Lemma 4.5, this is explicit in its proof. This then carries over to Theorem 3.5 if in the proof of Proposition 3.3 one lets $P_{0}$ run over sets of Dirichlet density 0 rather than only over finite sets. It also carries over to Theorem 4.6 (and similarly the theorems in the following two sections) if in the proof of Proposition 4.4 one lets $\mathcal{U}$ run only over those ultrafilters on $\mathbb{P}_{K}$ that contain the filter of subsets of Dirichlet density $1$ rather than over all nonprincipal ultrafilters.
@@ 5. Existential $\mathfrak{L}_{\rm ring}$-theory with parameters
If $(F,v)$ and $(F^{\prime},v^{\prime})$ are henselian nontrivially valued separable extensions of a trivially valued field $C$ such that $Fv$ and $F^{\prime}v^{\prime}$ have the same universal/existential $\mathfrak{L}_{\rm ring}(C)$-theory and every geometrically integral smooth affine $C$-variety has a rational point over both $Fv$ and $F^{\prime}v^{\prime}$, then $F$ and $F^{\prime}$ have the same universal/existential $\mathfrak{L}_{\rm ring}(C)$-theory, i.e. ($\ast$2.1) holds for $T=T_{3}$ with the weaker conclusion for $\mathfrak{L}_{\rm ring}(C)$ instead of $\mathfrak{L}_{\rm val}(C)$.
@@
We claim that by enlarging $C$ we can assume without loss of generality that $C$ is separably closed in $Fv$ and $F^{\prime}v^{\prime}$. Indeed, let $D$ be the separable algebraic closure of $C$ in $Fv$. Since $Fv$ and $F^{\prime}v^{\prime}$ have the same universal/existential $\mathfrak{L}_{\rm ring}(K)$-theory, the relative algebraic closure of $C$ in $Fv$ is isomorphic to the one in $F^{\prime}v^{\prime}$ [FJ08, Lemma 20.6.3(b)], and so we can choose an isomorphism of $D$ with the separable algebraic closure of $C$ in $F^{\prime}v^{\prime}$. As $v$ is trivial on $C$, we can identify the separable algebraic closure of $C$ in $F$ with a subfield of $D$ via the residue map, and since $v$ is henselian, this subfield is actually all of $D$ (cf. [AF16, Lemma 2.3]). Similarly we identify $D$ with the separable algebraic closure of $C$ in $F^{\prime}$. Note that $F$ and $F^{\prime}$ remain separable over $D$. Finally, every geometrically integral smooth affine $D$-variety has a rational point over $Fv$ and $F^{\prime}v^{\prime}$: Such a $D$-variety is the base change of a geometrically integral smooth affine $C^{\prime}$-variety $V$ for a finite subextension $C^{\prime}/C$ of $D/C$. The Weil restriction $W={\rm res}_{C^{\prime}/C}(V)$ is a geometrically integral smooth affine $C$-variety (see [Sch94, Prop. 4.9] and [Die01, Lemma 1.10]), hence $W(Fv)\neq\emptyset$ by assumption. From the morphism $W_{C^{\prime}}\rightarrow V$ (see [Sch94, 4.2.5]) we conclude that also $V(Fv)\neq\emptyset$, and similarly $V(F^{\prime}v^{\prime})\neq\emptyset$.
So from now on assume that $C$ is separably closed in $Fv$ and $F^{\prime}v^{\prime}$. We already argued that then it is also separably closed in $F$. Since $F/C$ is in addition separable, it is therefore regular. Similarly, $F^{\prime}/C$ is regular. Therefore, by [Dit25, Lemma 4.2], $F$ and $F^{\prime}$ have the same existential $\mathfrak{L}_{\rm ring}(C)$-theory, and therefore the same universal/existential $\mathfrak{L}_{\rm ring}(C)$-theory, as soon as every geometrically integral smooth affine $C$-variety $V$ has an $F$-rational point if and only if it has an $F^{\prime}$-rational point. But every such $V$ has an $Fv$-rational point by assumption, and this $Fv$-rational point lifts to an $F$-rational point of $V$ since $v$ is henselian (see for instance [BLR90, §2.3 Prop. 5]), and similarly to an $F^{\prime}$-rational point. This concludes the proof. ∎
@@
[…]
@@
where now $\forall/\exists$ denotes universal/existential $\mathfrak{L}_{\rm ring}(K)$-consequences. As $\bigcap_{\mathcal{U}}{\rm Th}(K_{\mathcal{U}}v_{\mathcal{U}})=\bigcap_{\mathcal{U}}{\rm Th}(\prod_{v\in\mathbb{P}_{K}}Kv/\mathcal{U})$ is the $\mathfrak{L}_{\rm ring}(K)$-theory of almost all $Kv$, which by Lemma 4.5 is the $\mathfrak{L}_{\rm ring}(K)$-theory of pseudofinite fields containing $K$, the first claim follows. The decidability follows from Lemma 2.1(b) and Lemma 4.5. ∎